OpenStudy (anonymous):

For the points (1, -4) and (-2, -10), find the slope-intercept equation of the line passing through the points. I know the equation is y=mx+b and that m=2, but I don't know which numbers to plug in and where.

OpenStudy (callisto):

\[\frac{y - y_1}{x-x_1} = \frac{y_2-y_1}{x_2-x_1}\] Put \((x_1, y_1) = (1, -4) \) and \((x_2, y_2) = (-2, -10)\) into the formula. \[\frac{y - (-4)}{x-(1)} = \frac{-10-(-4)}{(-2)-(1)}\]\[\frac{y+4}{x-1} = \frac{-6}{-3}\]\[\frac{y+4}{x-1} = 2\]\[y+4 = 2(x-1)\] Now simplify it. Special note: As you can see, the right side of the formula is actually finding the slope of the line :)

OpenStudy (anonymous):

This is pretty confusing >< Sorry.

OpenStudy (callisto):

Hmm?! Which part?? No worries btw :)

OpenStudy (anonymous):

My professor taught it a different way, I think. I just can't find my notes.

OpenStudy (anonymous):

I remember it was a linear equation (no fractions)

OpenStudy (callisto):

\(y - y_1 = m (x-x_1)\)???

OpenStudy (anonymous):

Yeah, that was it! I just found my notes :P (I see that your way works, too. But, I have a final tomorrow. I don't want to have to learn something and unlearn something else last minute. Again, sorry)

OpenStudy (callisto):

never mind :) It's just the same though :P How do you find m generally with 2 points \((x_1, y_1)\) and \((x_2, y_2)\) given?

OpenStudy (anonymous):

The rise/run equation. m= (y2-y1)/(x2-x1)

OpenStudy (callisto):

Yes. So, put m = (y2-y1)/(x2-x1) into y-y1 = m (x-x1) You'll get \[y-y_1 = \frac{(y_2-y_1)}{(x_2-x_1)} (x-x_1)\]Now divide both sides by (x-x1), what will you get?

OpenStudy (anonymous):

Is the final answer y=2x-6?

OpenStudy (callisto):

Yes, I think. Actually, I was trying to tell you that it's not something new. You just have to combine what you've learn and change a little to get that. Anyway, use the method you're comfortable at - This is the best for you when you do your exam. Good luck!

OpenStudy (anonymous):

Thank you so much for your help! :)

OpenStudy (callisto):

Welcome :)